# Practical Insights into Knowledge Distillation for Pre-Trained Models

### v1

@@ Abstract
This research investigates the enhancement of knowledge distillation (KD) processes in pre-trained models, an emerging field in knowledge transfer with significant implications for distributed training and federated learning environments. These environments benefit from reduced communication demands and accommodate various model architectures. Despite the adoption of numerous KD approaches for transferring knowledge among pre-trained models, a comprehensive understanding of KD’s application in these scenarios is lacking. Our study conducts an extensive comparison of multiple KD techniques, including standard KD, tuned KD (via optimized temperature and weight parameters), deep mutual learning, and data partitioning KD. We assess these methods across various data distribution strategies to identify the most effective contexts for each. Through detailed examination of hyperparameter tuning, informed by extensive grid search evaluations, we pinpoint when adjustments are crucial to enhance model performance. This paper sheds light on optimal hyperparameter settings for distinct data partitioning scenarios and investigates KD’s role in improving federated learning by minimizing communication rounds and expediting the training process. By filling a notable void in current research, our findings serve as a practical framework for leveraging KD in pre-trained models within collaborative and federated learning frameworks.
@@ I Introduction
Knowledge distillation (KD), initially conceptualized for model compression [1, 2, 3], has evolved to facilitate knowledge transfer between pre-trained models[4, 5, 6, 7]. Notably, these pre-trained models might be exposed to distinct data samples, resulting in performance variations due to various heterogeneities. One of these is statistical heterogeneity, where models are trained with data that are not independent and identically distributed (non-IID). Additionally, system heterogeneity introduces challenges, as devices vary in their available bandwidth and computational capabilities. Learning across such diverse systems and imbalanced datasets is complex yet critical.
The application of KD in distributed training and federated learning environments is increasingly recognized for its potential to mitigate communication overhead and facilitate diverse model architectures [6, 7, 8, 4, 5, 9, 10, 11]. Despite this growing interest, there remains a notable gap in understanding KD’s application, particularly in pre-trained models. Recent studies employ various KD algorithms—such as vanilla KD, deep mutual learning [4, 9], and data partitioning KD[12]—to facilitate knowledge transfer between pre-trained models. These methods’ effectiveness across varied data distribution strategies and their comparative advantages or limitations have not been systematically investigated.
Additionally, the realm of hyperparameter optimization, particularly concerning parameters such as temperature and weight, remains largely unexplored, presenting a significant opportunity for advancement in this domain [13].
This paper aims to address these gaps by conducting an in-depth study on joint distillation [13], a scenario where distributed participants engage in direct peer-to-peer knowledge exchange between pre-trained models. We present a comparative analysis of diverse KD approaches, considering various data partitioning strategies. Our exploration includes an assessment of the influence of transfer set options and a detailed investigation into hyperparameter optimization. Through extensive grid search results, we identify practical guidelines for tuning to enhance model performance. This study aims to enrich the understanding of KD techniques in pre-trained models and provide actionable insights for optimizing these methods in distributed and federated learning contexts.
This paper advances the field of knowledge distillation in distributed training systems with the following contributions:
  • We conduct a comprehensive comparative analysis of KD approaches across five data partitioning strategies and assess the impact of different transfer set options.
  • We examine hyperparameter optimization, focusing on temperature and weight, using grid search to ascertain when tuning is most beneficial and to determine optimal tuning scenarios for diverse data partitions.
@@
  • We demonstrate the efficiency of KD in federated learning (FL), evidenced by reduced training rounds to reach desired accuracy levels.
We plan to openly release the source code associated with this study.
II Preliminaries and related work
II-A Knowledge distillation
KD is a technique initially proposed to transfer knowledge from a large, complex model (termed the teacher) to a smaller model (termed the student), with the goal of maintaining comparable performance [1, 2, 3]. The central concept is that the student model can learn more effectively by emulating the soft targets, which are the output probabilities computed using the softmax function parameterized by temperature $T$, as shown:
@@ II-A Knowledge distillation
Various KD methodologies utilize different loss functions and datasets, often referred to as a transfer set or proxy dataset, which can include public labeled or unlabeled datasets, or the teacher’s training data. When labeled data is available, the student model’s training combines the cross-entropy loss with the teacher’s soft targets and a distillation loss as:
@@
II-B Knowledge distillation with pre-trained models
Beyond model compression, KD’s applicability includes scenarios where both the teacher and student models are pre-trained on different data subsets. This application of KD aims to consolidate the distinct knowledge each model possesses, a concept increasingly relevant in distributed training and federated learning [4, 5, 9, 6, 11, 10, 7, 12].
Our study focuses in the following approaches:
@@ II-B Knowledge distillation with pre-trained models
Tuned KD
Tuned KD refers to the optimization of KD hyperparameters, particularly the temperature ($T$) and weight ($\alpha$). This approach adjusts the standard KD framework to better suit the characteristics and learning capacity of the pre-trained student model.
@@
In this context, $m^{t}_{l}$ and $m^{st}_{l}$ are binary masks for each sample $l$, determining whether knowledge transfer from the teacher model $t$ is appropriate, or if the student model $s$ should retain its own knowledge. The distillation loss function is then defined as:
@@
In the unsupervised variant of DP-KD, samples are assigned based on the maximum prediction probability for a sample, selecting the model with the highest confidence in its prediction. This model selection by confidence circumvents the need for true labels, enabling knowledge transfer to occur even without labeled data. We explore the efficacy of both the supervised and unsupervised paradigms of DP-KD to assess their respective impacts on the performance of pre-trained models.
II-C Related work
KD has seen notable expansion into distributed training, particularly in the transfer of knowledge among pre-trained models. This expansion has been increasingly prominent in fields such as distributed training and FL, as evidenced by a range of studies [4, 5, 9, 6, 11, 10, 7]. Techniques including FedMD [5], Cronus [11], and DS-FL [6] leverage KD for enhanced communication efficiency and to address model heterogeneity challenges. Decentralized frameworks like Def-KT[4] and ProxyFL [9] employ online distillation, facilitating collaborative learning in the absence of a central authority.
Despite these advancements, much of the research remains focused on singular KD strategies and a narrow range of data partitioning methods. Our study seeks to fill this gap by examining a variety of KD methodologies across different data distribution and transfer set scenarios, aiming to detail the performance nuances of diverse KD techniques in distributed training. This includes exploring the impact of hyperparameter tuning in KD, as highlighted by recent findings [13]. Furthermore, while existing FL studies have successfully integrated KD to reduce communication loads [6, 7, 8, 4], and support flexible model architectures across participants [5, 9, 10, 11], our study proposes a new approach to further cut communication overhead, particularly in cross-silo FL settings, enhancing FL system efficiency.
@@ III Methodology
Based on the assumption of a set of participants (so-called clients), each with a private dataset, this study focuses on the application of KD techniques between pre-trained models, performed as follows:
  1. [itemsep=0pt, parsep=0pt]
@@
    By default, the student model’s training dataset is the transfer set. We also examine the use of public datasets, both labeled and unlabeled, to evaluate their impact on the distillation process.
@@
Task, models, and datasets.  For this study, we focus on a standard computer vision task: image classification. As a dataset, we use CIFAR10 [16], which is commonly used in the FL literature (e.g., [17, 18, 19, 20, 21, 22]). The CIFAR-10 dataset consists of 32x32 color images of 10 classes. There are 50K training images and 10K test images. We use 10% of the training set as a validation set. CIFAR-10’s testing set is used as a shared dataset to measure the accuracy of the participants’ models before and after distillation.
To exclude the effects of different model architectures, we use ResNet-18 for each participant. We plan to revisit this choice in follow-up work.
For each participant, we pre-train the model over its training dataset for at most 100 epochs. We employ early stopping after the validation performance plateaus for 10 epochs. We use the Adam optimizer with a learning rate of $1\times 10^{-3}$, weight decay of $4\times 10^{-4}$, and a batch size of 32. These hyperparameters remain constant across all experiments.
@@
In the Tuned KD approach, we primarily focus on the optimization of the $\alpha$ parameter, followed by the tuning of the temperature parameter $T$. This optimization is conducted via a systematic grid search across the parameter space, as delineated in Table I.
@@
For simplicity and to align with our peer-to-peer collaborative system assumption, we predominantly focus on a single student and teacher setting. While we briefly delve into configurations with multiple teachers and a single student in § V, we acknowledge the complexity of these setups and reserve their extensive exploration for future research endeavors.
@@
Also, to obtain a balanced perspective on both the gains and potential drawbacks of KD, we assess forgetting, quantified as the decrease in per-class accuracy of the student model post-distillation relative to its pre-distillation performance, and learning, which is the increase in per-class accuracy of the student model post-distillation.
IV Multidimensional Evaluation of Knowledge Distillation Techniques
@@
We start with a comparison of various KD approaches, demonstrating their performance across different data distribution scenarios and transfer set configurations. Fig. 2 illustrates the results. Each subplot of the figure is a scatter plot showing knowledge transfer across all 90 participant pairs. That is, each data point in these subplots corresponds to a distinct knowledge transfer instance between two models. The x-axis denotes the performance differential between the teacher and the student models (quantified as teacher accuracy minus student accuracy), while the y-axis value reflects the accuracy gain achieved through KD.
Teachers that exhibit notably higher accuracy compared to their corresponding students are categorized as strong, while those with lower accuracy relative to the students are designated as weak. Instances of KD that result in improved accuracy are designated as positive knowledge transfers.
This comparative analysis includes transfer sets of varying sizes to determine their impact on the efficacy of KD. Specifically, the “Public unlabeled (large)” transfer set comprises 5,000 samples, the “Public unlabeled (small)” and the “Public labeled transfer set” each contains 500 samples, randomly sampled from the original dataset before partitioning. The “Student data as a transfer set” varies in size corresponding to the student dataset.
We now analyze our main findings from this comparison. We first contrast the performance of DML and DP-KD with Vanilla KD, which represents the baseline approach. Then, we illustrate the impact of hyperparameter tuning by comparing Tuned KD against other approaches.
@@ IV-A Different transfer set options
Our analysis reveals a proportional relationship between the size of the transfer set and KD effectiveness. An increase in the transfer set’s sample size consistently leads to enhanced outcomes of positive knowledge transfers, as illustrated in Fig. 2 (particularly evident when comparing positive knowledge transfers in columns 2 and 3). These findings underscore the importance of transfer set size in optimizing KD efficiency and offer practical insights for refining KD methodologies. Prior work [27] supports this observation, suggesting that an expanded transfer set can significantly reduce the student classifier’s loss, aligning the student’s parameters more closely with those of the teacher model.
It is worth noting that a larger sample size in the transfer set may also lead to more pronounced negative knowledge transfers under certain conditions, as observed with the DML approach in Fig. 2 (compare columns 2 and 3). As will be discussed next, the DML approach is not recommended in the absence of a student dataset or another public labeled transfer set.
Finding (2): The DML approach exhibits poor performance in unsupervised transfer set settings.
Our findings indicate that while DML generally outperforms Vanilla KD and DP-KD in scenarios involving weaker teachers, particularly with ‘Student data as a transfer set’ and ‘Public labeled transfer set’, its performance significantly diminishes in the absence of labeled data. In such cases, DML often results in negative knowledge transfers, as observed in columns 2 and 3 of Fig. 2. This pattern cautions against employing DML when neither student data nor a labeled transfer set is accessible.
@@
Fig. 2 provides a detailed comparative analysis underscoring this finding. In scenarios such as Uniform Data Distribution, Quantity Skew Distribution, and Label Skew Random Chunks Distribution, Vanilla KD’s performance is comparable to that of DP-KD. However, it notably excels in the Label Skew Dirichlet Distribution scenario and demonstrates significant superiority in the Specialized Non-IID Distribution context, across all tested transfer set options. This pattern of performance highlights the robustness and adaptability of Vanilla KD, particularly in complex data distribution scenarios and when paired with teachers of comparable or superior strength.
@@
This outcome suggests that the model’s guidance may be less effective in scenarios characterized by significant class imbalances—even though the teacher model has been exposed to samples from all classes. Such a scenario can lead to misleading confidence from the teacher model, ultimately resulting in suboptimal knowledge transfer and poor performance outcomes. This underscores the importance of considering class balance in the teacher’s training dataset for effective DP-KD application.
@@
In contrast to Vanilla KD and DP-KD, DML exhibits a distinctive ability to facilitate positive knowledge transfers with weaker teachers when utilizing student data as the transfer set. This is clearly demonstrated in Fig. 2, particularly in column 1. Such findings highlight DML’s potential in scenarios where the teacher model is less robust, emphasizing its effectiveness in optimally leveraging student data. However, in scenarios involving stronger teachers, Vanilla KD tends to outperform DML.

### v2

# Analysis and Tuning of Knowledge Distillation for Efficient Collaborative Learning

@@ Abstract
Collaborative learning addresses key challenges in distributed systems, such as large-scale training, data privacy, and computational heterogeneity, by enabling decentralized entities to collaborate without centralizing data. This distributed computing paradigm is particularly crucial for heterogeneous environments, where maintaining communication efficiency and supporting diverse model architectures are vital. Among the various approaches to collaborative learning, knowledge distillation (KD) has emerged as a promising method, facilitating efficient knowledge transfer while minimizing communication overhead. However, existing research on KD often lacks a systematic empirical analysis of its effectiveness across distributed and heterogeneous settings.
This paper bridges this gap by conducting a comprehensive experimental analysis of KD methods for distributed systems. We evaluate vanilla KD, tuned KD, deep mutual learning, and data-partitioning KD across diverse data distributions and system configurations. Our results demonstrate that tuning KD hyperparameters, such as temperature and weight, enhances accuracy gains by 16% in uniform distributions and achieves an average improvement of 139% across non-IID scenarios compared to the best non-tuned approach. Furthermore, we propose a novel solution which leverages KD to enhance training efficiency in centralized collaborative learning and FL settings. By consolidating teacher models into a unified starting point, this approach achieves an average reduction of 60.6% in total communication rounds, effectively accelerating the training process. These findings provide insights and practical guidelines for collaborative learning in distributed systems.
Index Terms:
Knowledge Distillation, Collaborative Learning, Data Heterogeneity, Hyperparameter Tuning
@@ I Introduction
Knowledge Distillation, initially conceptualized for model compression [1, 2, 3], has evolved to facilitate knowledge transfer between pre-trained models[4, 5, 6, 7]. Notably, these pre-trained models might be exposed to distinct (i.e., unseen or out of distribution) data samples, resulting in performance variations due to various types of heterogeneity. One of these is statistical heterogeneity, where models are trained with data that are not independent and identically distributed (non-IID). Additionally, system heterogeneity introduces challenges as devices vary in their available bandwidth and computational capabilities. Learning across such diverse systems and imbalanced datasets is complex yet critical.
The application of KD in decentralized collaborative learning and federated learning environments is increasingly recognized for its potential to mitigate communication overhead and facilitate diverse model architectures [6, 7, 8, 4, 5, 9, 10, 11, 12, 13]. Despite this growing interest, a notable gap exists in understanding KD’s application, particularly in pre-trained models. Recent studies employ various KD algorithms—such as vanilla KD, deep mutual learning [4, 9], and data partitioning KD [14]—to facilitate knowledge transfer between pre-trained models. These methods’ effectiveness across varied data distribution strategies and their comparative advantages or limitations have not been systematically and empirically investigated. Additionally, KD hyperparameters tuning, particularly concerning parameters such as temperature and weight, remains largely unexplored, presenting a significant opportunity for advancement in this domain [15].
This paper addresses these gaps by conducting an in-depth study on joint (or collaborative) distillation. In this scenario, distributed participants use their pre-trained models to engage in direct peer-to-peer knowledge exchange. We present a comparative analysis of diverse KD approaches, considering various data partitioning strategies. Our exploration includes an assessment of the influence of transfer set options and a detailed investigation into hyperparameter optimization. Through extensive grid search results, we identify practical guidelines for tuning to enhance model performance. This study aims to enrich the understanding of KD techniques in pre-trained models and provide actionable insights for optimizing these methods in distributed and collaborative learning contexts.
This paper makes the following contributions to the field of knowledge distillation for collaborative learning systems:
  • We conduct a comprehensive comparative analysis of KD approaches across five data partitioning strategies and assess the impact of different knowledge transfer set methods and hyperparameters settings. [^1] We plan to release the source code associated with this study openly.
  • We examine hyperparameters optimization, focusing on temperature and weight, using grid search to ascertain when tuning is most beneficial and to determine optimal tuning scenarios for diverse data partitions.
@@
  • We explore the criteria for effective teacher model selection across data scenarios.
  • We demonstrate the efficiency of KD in the centralized collaborative learning and FL settings, evidenced by reduced training rounds to reach desired accuracy levels.
II Background and Related Work
This section introduces the concept of collaborative learning, elaborates on the foundations of KD, and discusses related work in these fields.
II-A Collaborative Learning
Collaborative learning is a paradigm where multiple models or agents work together to enhance collective performance through shared learning experiences. This model allows each device or node to learn independently while also enabling selective collaboration by transferring knowledge when necessary [16]. Collaborative learning encompasses both centralized approaches, such as FL [17], and decentralized peer-to-peer approaches [16, 18, 19, 11, 14].
Collaborative learning addresses critical challenges such as data privacy, bandwidth limitations, and data locality. However, it faces significant hurdles due to system and data heterogeneity [20, 21, 22, 23, 24, 25]. System heterogeneity arises from variations in hardware capabilities, network conditions, and available computational power across devices. This disparity can make it challenging to fit the same model across all participating devices uniformly and often leads to inefficient training processes, compromising the learning performance. Additionally, data heterogeneity—characterized by the non-IID nature of data distribution across different nodes—further complicates the scenario by producing models that do not generalize well across varied data distributions [26, 17, 12].
II-B Knowledge Distillation
KD is a technique initially proposed to transfer knowledge from a large, complex model (termed the teacher) to a smaller model (termed the student) to maintain comparable performance [1, 2, 3]. The central concept is that the student model can learn more effectively by emulating the soft targets, which are the output probabilities computed using the softmax function parameterized by temperature $T$, as shown:
@@ II-B Knowledge Distillation
Various KD methodologies utilize different loss functions and datasets, often referred to as a transfer set or proxy dataset, which can include publicly labelled or unlabeled datasets or the teacher’s training data. When labelled data is available, the student model’s training combines the cross-entropy loss with the teacher’s soft targets and a distillation loss as:
@@
II-C Knowledge Distillation With Pre-trained Models
Beyond model compression, KD’s applicability includes scenarios where both the teacher and student models are pre-trained on different data subsets. This application of KD aims to consolidate the distinct knowledge each model possesses, a concept increasingly relevant in decentralized collaborative learning  [11, 7, 14] and federated learning [4, 5, 9, 6, 14].
Therefore, our study focuses on the analysis of the following knowledge distillation (KD) approaches:
@@ II-C Knowledge Distillation With Pre-trained Models
In this context, $m^{t}_{l}$ and $m^{st}_{l}$ are binary masks for each sample $l$, determining whether knowledge transfer from the teacher model $t$ is appropriate, or if the student model $s$ should retain its own knowledge. The distillation loss is defined as:
@@
In the unsupervised variant of DP-KD, samples are assigned based on the maximum prediction probability for a sample, selecting the model with the highest confidence in its prediction. This model selection by confidence circumvents the need for true labels, enabling knowledge transfer even without labeled data. We explore the efficacy of both the supervised and unsupervised paradigms of DP-KD to assess their respective impacts on the performance of pre-trained models.
II-D Related work
To address the limitations of system and data heterogeneity in collaborative learning and reduce communication overhead, KD has seen notable expansion, particularly in knowledge transfer among pre-trained models, as evidenced by a range of studies [4, 9, 10, 7]. Techniques such as FedMD [5], Cronus [11], and DS-FL [6] leverage KD to improve communication efficiency and address model heterogeneity challenges. Furthermore, decentralized learning frameworks like Def-KT [4] and ProxyFL [9] employ online distillation to facilitate collaborative learning without a central authority.
Despite these advancements, much of the research remains focused on singular KD strategies and a narrow range of data partitioning methods. Our study seeks to fill this gap by examining a variety of KD methodologies across different data distribution and transfer set scenarios, aiming to detail the performance nuances of diverse KD techniques in distributed training. This includes exploring the impact of hyperparameter tuning in KD, as highlighted by recent findings [15]. Furthermore, while existing FL studies have successfully integrated KD to reduce communication loads [6, 7, 8, 4] and support flexible model architectures across participants [5, 9, 10, 11], our study proposes a new approach to further cut communication overhead, particularly in cross-silo FL settings, enhancing FL system efficiency.
@@ III Methodology
Overview.  Our methodology is structured to evaluate KD in collaborative learning environments. This research systematically compares different KD approaches under diverse scenarios, aiming to enhance our understanding of KD’s application between pre-trained models, pinpoint the most effective KD strategies for specific conditions, and address existing gaps in the literature. Our experimental framework begins with a foundational study in a controlled peer-to-peer setting involving a direct KD interaction between one student and one teacher. Subsequently, (in IV-G), we extend the scope to include scenarios involving multiple teachers, simulating more complex, centralized collaborative learning settings and federated learning environments.
III-A Proposed Approach
We aim to advance KD for collaborative learning by analyzing under-explored aspects and proposing novel solutions to address efficiency challenges in distributed settings. Our focus includes three key techniques: Tuned KD, Model Amalgamation, and Adaptive Teacher Weighting, each evaluated to provide actionable insights. Additionally, we introduce a novel pre-consolidation method for federated learning, which is discussed in later sections.
Tuned KD.  In this work, we study the tuning of KD hyper-parameters. Hence, our proposed approach, Tuned KD, refers to the optimization of KD hyperparameters, particularly the temperature ($T$) and weight ($\alpha$) hyper-parameters. This approach adjusts the standard KD framework to fit better the characteristics and learning capacity of the pre-trained student model. We aim to leverage this approach to systematically assess KD’s efficacy and adaptability across different collaborative configurations.
Model Amalgamation.  We also study combining multiple models to amalgamate knowledge across multiple pre-trained models via KD effectively (which is analogous to ensemble learning). Our focus includes critical aspects such as selecting an appropriate starting student model and assessing various teacher weighting strategies across different transfer set scenarios.
Weighing Teachers.  Traditionally, equal weighting is used for all teachers; we investigate adaptive weighting, which is based on each teacher’s class-wise contribution. This weighting mechanism was inspired by a method outlined in the study [29], adapted to use per-class accuracy instead of per-class label counts to ensure generalizability when label count information is unavailable.
Here, $J$ is the total number of teachers, and the dimension of each $\alpha_{j}$ vector matches the number of classes. The $\mathcal{L}_{KL}$ in Eq. 2 can be updated considering the combined influence of multiple teachers, as:
Figure 2: Pre-Consolidation of Models using KD. Multiple teacher models contribute their knowledge to create a consolidated model, which is used as the initial global model for the FL process.
III-B Evaluation Method
Experimental Setup.  Based on the assumption of a set of participants (so-called clients), each with a private dataset, this study focuses on the application of KD techniques between pre-trained models, performed as follows:
@@ III-B Evaluation Method
    By default, the student model’s training dataset is the transfer set. We also examine the use of public datasets, both labelled and unlabeled, to evaluate their impact on the distillation process.
@@
Task, models, and datasets.  This study focuses on a standard computer vision task, i.e., image classification. As a dataset, we use CIFAR-10 [30] and CINIC-10 [31], which are commonly used in the FL literature (e.g., [32, 33, 34, 35, 36, 37, 38, 39]). The CIFAR-10 dataset consists of 32x32 color images of 10 classes. There are 50K training images and 10K test images. We use 10% of the training set as a validation set. The CINIC-10 dataset comprises 270,000 images divided into training, validation, and test sets, each containing 90,000 images across 10 classes. We merge the training and validation sets, utilizing 1% of this set for validation and the remaining 99% for training. The testing set of each dataset is used to measure the accuracy of the participants’ models before and after distillation. CIFAR-10 is the default data set, unless otherwise specified.
We utilize the ResNet-18 architecture [40] for all participants to assess the effects of knowledge distillation techniques and data strategies employed, thereby eliminating confounding variables introduced by differing architectures.
We pre-train the model over its training dataset for at most 100 epochs for each participant. We employ early stopping after the validation performance plateaus for 10 epochs. We use the Adam optimizer with a learning rate of $1\times 10^{-3}$, weight decay of $4\times 10^{-4}$, and a batch size 32. These hyperparameters remain constant across all experiments.
@@
In the Tuned KD approach, we primarily focus on optimizing the $\alpha$ parameter, followed by tuning the temperature parameter $T$. This optimization is conducted via a systematic grid search across the parameter space, as delineated in Table I.
@@
In our analysis, we predominantly focus on a single student and teacher setting, aligning with the assumption of a peer-to-peer collaborative system. Following this, in § IV-G, we briefly explore configurations involving multiple teachers and a single student, recognizing their potential to significantly impact centralized collaborative learning systems. We, however, acknowledge the complexity of these multi-teacher setups and reserve their extensive exploration for future research.
@@
IV Evaluation Results
@@
We compare various KD approaches and demonstrate their performance across different data distribution scenarios and transfer set configurations. Fig. 3 illustrates the results. Each subplot of the figure is a scatter plot showing knowledge transfer across all 90 participant pairs. Each data point in these subplots corresponds to a distinct knowledge transfer instance between two models. The x-axis denotes the performance difference between the teacher and the student models (quantified as teacher accuracy minus student accuracy), while the y-axis value reflects the accuracy gain achieved via KD.
Teachers with notably higher accuracy than their corresponding students are categorized as strong, while those with lower accuracy relative to the students are designated as weak. Instances of KD that result in improved accuracy are designated as positive knowledge transfers.
This comparative analysis includes transfer sets of varying sizes to determine their impact on the efficacy of KD. Specifically, the “Public unlabeled (large)” transfer set comprises 5,000 samples, and the “Public unlabeled (small)” and the “Public labeled transfer set” each contain 500 samples, randomly sampled from the original dataset before partitioning. The “Student data as a transfer set” varies in size, corresponding to the student dataset.
We now analyze our main findings from this comparison. First, we contrast the performance of DML and DP-KD with that of Vanilla KD, representing the baseline approach. Then, we illustrate the impact of hyperparameter tuning by comparing Tuned KD against other approaches.
@@ IV-A Different transfer set options
Our analysis reveals a proportional relationship between the size of the transfer set and the effectiveness of knowledge distillation. As demonstrated in Fig. 3, expanding the transfer set size consistently improves the outcomes of knowledge transfer. This improvement is particularly noticeable when comparing positive knowledge transfers in columns 2 and 3 of the figure. This enhancement in KD effectiveness with larger transfer sets can be attributed to the increased diversity and representativeness of the sample distribution. A larger set provides more comprehensive coverage of the input space, allowing the student model to learn a more generalized data representation. This principle aligns with statistical learning theory, which posits that larger datasets can reduce the variance of estimators and improve model generalization on unseen data [45]. Furthermore, supporting research [46] indicates that an expanded transfer set can significantly diminish the student classifier’s loss, thereby aligning the student’s parameters more closely with the teacher model.
It is worth noting that a larger sample size in the transfer set may also lead to more pronounced negative knowledge transfers under certain conditions, as observed with the DML approach in Fig. 3 (compare columns 2 and 3). As discussed next, the DML approach is not recommended without a student dataset or another public labeled transfer set.
Finding (2): The DML approach performs poorly in unsupervised transfer set settings.
Our findings indicate that DML generally outperforms Vanilla KD and DP-KD in scenarios involving weaker teachers, particularly in scenarios where ‘Student data’ and ‘Public labelled data’ are used as transfer sets. On the other hand, in the absence of labeled data, its performance significantly diminishes. In such unsupervised settings, DML can reinforce existing biases or errors between the models, as there is no ground truth to guide the correction of predictions. This issue is further exacerbated by the peer-to-peer learning structure of DML, where two models continuously adjust their parameters based on potentially flawed outputs from each other, leading to compounded inaccuracies without proper error correction mechanisms. This often results in negative knowledge transfers, as observed in columns 2 and 3 of Fig. 3. This pattern cautions against employing DML when neither student data nor a labeled transfer set is accessible.
@@
Fig. 3 provides a detailed comparative analysis underscoring this finding. In Uniform Data Distribution, Quantity Skew Distribution, and Label Skew Random Chunks Distribution scenarios, Vanilla KD’s performance is comparable to that of DP-KD. However, it notably excels in the Label Skew Dirichlet Distribution scenario and demonstrates significant superiority in the Specialized Non-IID Distribution context, across all tested transfer set options. This pattern of performance highlights the robustness and adaptability of Vanilla KD, particularly in complex data distribution scenarios and when paired with teachers of comparable or superior strength.
@@
DP-KD’s underperformance in the Specialized Non-IID Distribution scenario, which features significant class imbalances, highlights its challenges in handling skewed data distributions. Here, even though the teacher model has encountered samples from all classes, its potential overconfidence in over-represented classes can lead to misleading guidance for the student model. This misalignment often results in the reinforcement of errors rather than corrective learning. This underscores the importance of considering class balance in the teacher’s training dataset for effective DP-KD application.
@@
The limited effectiveness of Vanilla KD and DP-KD when paired with weaker teachers, particularly when using student data as the transfer set, underscores a significant drawback in these methods under specific circumstances. The performance decline can be theoretically linked to the weaker teachers’ insufficient guidance, which inadequately boosts the student model’s learning, especially when leveraging potentially less diverse student data
DML demonstrates an ability to extract useful knowledge from these interactions, even with weaker teacher models, as evidenced in Fig. 3, particularly in column 1. This indicates DML’s ability to mutually benefit participating models, effectively utilizing their collective insights despite individual limitations. However, Vanilla KD tends to surpass DML in scenarios involving stronger teachers. This indicates that the method choice is dependent on the teacher models’ strength, i.e., DML is preferred when teacher models are less powerful, and Vanilla KD is preferred when teacher models are robust.
IV-C Optimizing Hyperparameters: Tuned KD
Refer to caption Refer to caption        Refer to caption Refer to caption